Solving Problems Using Linear Systems
4. Rehman invests his summer earnings of $3050. He invests part of the money at 8%/year, and the rest at 7.5%/year. After 1 year, these investments earn $242 in simple interest. How much did he invest at each rate?
(answers are: $2650 at 8%/year and $400 at 7.5%/year, but i dont know how to solve it?)

Answer by stanbon(75887)

Rehman invests his summer earnings of $3050. He invests part of the money at 8%/year, and the rest at 7.5%/year. After 1 year, these investments earn $242 in simple interest. How much did he invest at each rate?
-----------------
Quantity Equation: x + y = 3050 dollars
Interest Equation:0.08x + 0.076y = 242
--------------------------------------------
Modify these equations so you can eliminate one of the variables:
80x + 80y = 244000
80x + 76y = 242000
------------------------
4y = 2000
y = 500 (amt invested at 7.6%
Substitute into x+y = 2050
x + 500 = 3050
x = 2550 (amt invested at 8%)
===================================
Cheers,
Stan H.